Equation of the straight line that passes through (0,-1) and is perpendicular to the line with equation 3x - 2y + 5 = 0
Answer by TimothyLamb(4379)

---
3x - 2y + 5 = 0
3x + 5 = 2y
y = (3/2)x + 5/2
---
perpendicular line has slope = -2/3 = -0.666666
---
given:
(0,-1)
---
to solve the linear equation, copy and paste this (the given point and the perpendicular slope):
0,-1,-0.666666
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (-1.5, 0)
y-intercept = (0, -1)
slope = -0.666666
---
answer:
the slope-intercept form of our new line:
y = (-0.666666)x - 1
y = (-2/3)x - 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php